OpenStudy (superhelp101):

How many milliliters of sodium metal, with a density of 0.97 g/mL, would be needed to produce 53.2 grams of hydrogen gas in the single-replacement reaction below? Show all steps of your calculation as well as the final answer. Na + H2O → NaOH + H2

OpenStudy (superhelp101):

@Abhisar can you check my explanation for this question?

OpenStudy (superhelp101):

Step: 1. Use the formula: Number of Moles=Mass(g)/Molar Mass(g/mol) 2. So then you have to plug in the numbers from the question to the formula, it should look like this 53.2/2=26.6 3. Then take a look at the Mole Ratio of Na:H2 4. The ratio is 2:1 5. x:26.6 6. Then you multiple to find the moles 26.6 X 2/1=53.2 moles 7. Next use this formula to solve for mass: Mass(g)=Molar Mass(g/mol) X Number of moles 8. Plug in the numbers to the equation from above to find the grams of Sodium 53.2*23.0=1223.6 grams of Sodium 9. Use this formula to determine the density: Density=Mass/Volume 10. Plug in the information from the question to the formula to find density: 1223.6/V=0.97 0.97V=1223.6 11. Then you come up with your answer!: V=1261.4 ml

OpenStudy (abhisar):

Did u balanced the equation first ?

OpenStudy (superhelp101):

no i forgot, is it 2 Na + 2 HOH --> 2NaOH + H2

OpenStudy (abhisar):

yep !

OpenStudy (superhelp101):

so do I do those steps after i balanced the equation

OpenStudy (abhisar):

Getting it ?

OpenStudy (superhelp101):

how do i figure out how many grams of sodium metal is required to react with 1 gram of hydrogen gasw many

OpenStudy (abhisar):

By comparing the ratios, i showed that in ur last question.

OpenStudy (superhelp101):

oh

OpenStudy (superhelp101):

let me go back to that

OpenStudy (abhisar):

Sorry there is an error in the steps i mentioned !

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\large \sf \ Step~1\) Find how many grams of sodium metal is required to produce 1 gram of hydrogen gas \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\large \sf \ Step~2\) Find how many grams of sodium metal is required to produce 53.2 grams of hydrogen. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\large \sf \ Step~3\) Convert the mass (grams) of sodium metal into volume (ml) by using given density. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

2 moles of Na are required to produce 1 mole of Hydrogen, i.e 45.97 grams of Na is required to produce 2 gram of hydrogen.

OpenStudy (abhisar):

so 22.98 grams of Na is required for producing 1 gram of hydrogen. Can u solve it futher ?

OpenStudy (abhisar):

*further

OpenStudy (superhelp101):

let me see

OpenStudy (superhelp101):

i am stuck

OpenStudy (abhisar):

Where ?

OpenStudy (superhelp101):

i just i don't know what do. should i set up and equation?

OpenStudy (abhisar):

but i explained u everything in ur last question...

OpenStudy (abhisar):

ok do one thing....rewrite the steps u know using the balanced equation.

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @superhelp101 Step: 1. Use the formula: Number of Moles=Mass(g)/Molar Mass(g/mol) 2. So then you have to plug in the numbers from the question to the formula, it should look like this 53.2/2=26.6 3. Then take a look at the Mole Ratio of Na:H2 4. The ratio is 2:1 5. x:26.6 6. Then you multiple to find the moles 26.6 X 2/1=53.2 moles 7. Next use this formula to solve for mass: Mass(g)=Molar Mass(g/mol) X Number of moles 8. Plug in the numbers to the equation from above to find the grams of Sodium 53.2*23.0=1223.6 grams of Sodium 9. Use this formula to determine the density: Density=Mass/Volume 10. Plug in the information from the question to the formula to find density: 1223.6/V=0.97 0.97V=1223.6 11. Then you come up with your answer!: V=1261.4 ml \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

correct these one

OpenStudy (abhisar):

brb..then i'll check

OpenStudy (superhelp101):

correct what? i want to go back to my last question but idk how, do i just go on closed questions :/

OpenStudy (superhelp101):

1. Na + H2O → NaOH + H2 in mol grams 2. 23 + 18 → 40 + 2 3. You need 23 gr of Na to produce 2 gr of H2, and you will need = 53.2*23/2 = 611.8 gr of Na to yield 53.2 gr of H2 4. 611.8 * 0.97 = 593.45 ml of Na is this right?

OpenStudy (superhelp101):

@Abmon98 what do i need to do?

OpenStudy (abmon98):

1. Na + H2O → NaOH + H2 in mol grams 2. 23 + 18 → 40 + 2 3. You need 23 gr of Na to produce 2 gr of H2, and you will need = 53.2*23/2 = 611.8 gr of Na to yield 53.2 gr of H2 4. 611.8 * 0.97 = 593.45 ml of Na Thats an incorrect method of solving you need to first find the number of moles or the amount that reacted to produce 53.2 moles of hydrogen not just from the atomic weight and then comparing ratios.

OpenStudy (superhelp101):

oh oops

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @superhelp101 1. 2Na + 2H2O → 2NaOH + H2 in mol grams 2. You need 46 gr of Na to produce 2 gr of H2, and you will need = 23 g of Na to produce 1 g of H2

OpenStudy (abmon98):

What you need to do is to solve to obtain the number of moles of hydrogen, use mole ratio from the balanced chemical equation. Use Number of moles=Mass(g)/Molar Mass(g/mol) to find out the mass of sodium consumed to produce 53.2 grams of hydrogen and use Density=mass/volume. Density is provided and mass too solve to get volume

OpenStudy (superhelp101):

@Abmon98 i see i will have to do that. @Abhisar do you agree?

OpenStudy (abhisar):

yeah obvsly ! :) check the steps i told u above..same thing !

OpenStudy (superhelp101):

ah. ok. thank you both!!!!!!!!!!!!! i wish i can best response you both :'(

OpenStudy (abhisar):

aah..it's ok :)

OpenStudy (abmon98):

Your welcome :D